OpenStudy (anonymous):

please show your work 12 is 30% of what number Please show your work

OpenStudy (anonymous):

12=(30)/(100)*x (30)/(100)*x=12 (3)/(10)*x=12 (3x)/(10)=12 3x=120 x=40